# Transport properties in a model of confined granular mixtures at moderate densities

### v1

@@ Acknowledgements.
We acknowledge financial support from Grant No. PID2024-156352NB-I00 and from Grant No. GR24022 funded by Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe”. The research of David González Méndez has been supported by the predoctoral fellowship FPU24/01056 from the Spanish Government. AUTHOR DECLARATIONS
@@ Appendix A First-order approximation
$\partial_{t}^{(0)}f_{i}^{(1)}-\sum_{j=1}^{s}\;J_{ij}^{(1)}[f_{i},f_{j}]=-\left(D_{t}^{(1)}+\mathbf{V}\cdot\nabla+\mathbf{g}\cdot\frac{\partial}{\partial\mathbf{v}}\right)f_{i}^{(0)},$ (135)
where $D_{t}^{(1)}=\partial_{t}^{(1)}+\mathbf{U}\cdot\nabla$ and the first order contribution $\sum_{j=1}^{s}\;J_{ij}^{(1)}[f_{i},f_{j}]$ to the Enskog collision operator is given by Garzó (2019)
@@
The balance equations at this order provide the actions of the operators $D_{t}^{(1)}$ over the hydrodynamic fields as
$D_{t}^{(1)}n_{i}=-n_{i}\nabla\cdot\mathbf{U},\quad D_{t}^{(1)}\mathbf{U}=-\rho^{-1}\nabla p+\mathbf{g},$ (139)
$\frac{d}{2}nD_{t}^{(1)}T=-p\nabla\cdot\mathbf{U}-\frac{d}{2}nT\zeta_{U}\nabla\cdot\mathbf{U},$ (140)
@@
$\left(D_{t}^{(1)}+\mathbf{V}\cdot\nabla+\mathbf{g}\cdot\frac{\partial}{\partial\mathbf{v}}\right)f_{i}^{(0)}=\sum_{j=1}^{s}\Bigg(\mathbf{V}n_{j}\frac{\partial f_{i}^{(0)}}{\partial n_{j}}+\frac{n_{j}}{\rho}\frac{\partial p}{\partial n_{j}}\frac{\partial f_{i}^{(0)}}{\partial\mathbf{V}}\Bigg)\cdot\nabla\ln n_{j}$
@@
$\partial_{t}^{(0)}f_{i}^{(1)}+\left({\cal L}f^{(1)}\right)_{i}=\mathbf{A}_{i}\cdot\nabla\ln T+\sum_{j=1}^{N}\mathbf{B}_{ij}\cdot\nabla\ln n_{j}+C_{i,\lambda\beta}\frac{1}{2}\left(\partial_{\beta}U_{\lambda}+\partial_{\lambda}U_{\beta}-\frac{2}{d}\delta_{\lambda\beta}\nabla\cdot\mathbf{U}\right)+D_{i}\nabla\cdot\mathbf{U},$ (142)
@@
$D_{i}^{\prime}\left(\mathbf{V}\right)=\frac{1}{d}\mathbf{V}\cdot\frac{\partial f_{i}^{(0)}}{\partial\mathbf{V}}+\left(\zeta^{(1,0)}+\frac{2}{d}\frac{p}{nT}\right)T\frac{\partial f_{i}^{(0)}}{\partial T}+\sum_{j=1}^{N}n_{j}\left\{\frac{\partial f_{i}^{(0)}}{\partial n_{j}}+\frac{1}{d}\mathcal{K}_{ij,\beta}\left[\frac{\partial f_{j}^{(0)}}{\partial V_{\beta}}\right]\right\}.$ (154)
@@ Appendix B Collisional contributions to the pressure tensor
$P_{\lambda\beta}^{(\Delta\neq 0)c}$ $=$ $2\sum_{i,j}\chi_{ij}m_{ij}\sigma_{ij}^{d}\Delta_{ij}\int d\mathbf{v}_{1}\int\mathrm{d}\mathbf{v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\,\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot\mathbf{g}_{12})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12})\widehat{\sigma}_{\lambda}\widehat{\sigma}_{\beta}\Bigg[f_{i}^{(1)}(\mathbf{V}_{1})f_{j}^{(0)}(\mathbf{V}_{2})$
$+\frac{1}{2}f_{i}^{(0)}(\mathbf{V}_{1})\frac{\partial f_{j}^{(0)}(\mathbf{V}_{2})}{\partial\mathbf{r}}\cdot{\boldsymbol{\sigma}}_{ij}\Bigg]=\frac{2B_{1}}{d+1}\sum_{i,j}\chi_{ij}m_{ij}\sigma_{ij}^{d}\Delta_{ij}$
$\times\int\mathrm{d}\mathbf{v}_{1}\int\mathrm{d}\mathbf{v}_{2}g_{12}^{-1}\left(g_{12,\lambda}g_{12,\beta}+g_{12}^{2}\delta_{\lambda\beta}\right)f_{i}^{(1)}(\mathbf{V}_{1})f_{j}^{(0)}(\mathbf{V}_{2})-\frac{B_{2}}{d+2}\sum_{i,j}\chi_{ij}m_{ij}\sigma_{ij}^{d+1}\Delta_{ij}\partial_{\nu}U_{\mu}$
$\times\int\mathrm{d}\mathbf{v}_{1}\int\mathrm{d}\mathbf{v}_{2}\left(g_{12,\nu}\delta_{\lambda\beta}+g_{12,\lambda}\delta_{\nu\beta}+g_{12,\beta}\delta_{\lambda\nu}\right)f_{i}^{(0)}(\mathbf{V}_{1})\frac{\partial f_{j}^{(0)}(\mathbf{V}_{2})}{\partial V_{2\mu}}=\frac{2B_{1}}{d+1}\sum_{i,j}\chi_{ij}m_{ij}\sigma_{ij}^{d}\Delta_{ij}$
$\times\int\mathrm{d}\mathbf{v}_{1}\int\mathrm{d}\mathbf{v}_{2}g_{12}^{-1}\left(g_{12,\lambda}g_{12,\beta}+g_{12}^{2}\delta_{\lambda\beta}\right)\Bigg[\mathcal{C}_{i,\mu\nu}(\mathbf{V}_{1})\frac{1}{2}\left(\partial_{\mu}U_{\nu}+\partial_{\nu}U_{\mu}-\frac{2}{d}\delta_{\mu\nu}\nabla\cdot\mathbf{U}\right)+\mathcal{D}_{i}\left(\mathbf{V}_{1}\right)\nabla\cdot\mathbf{U}\Bigg]$
$\times f_{j}^{(0)}(\mathbf{V}_{2})-\frac{B_{2}}{d+2}\sum_{i,j}\chi_{ij}m_{ij}n_{i}n_{j}\sigma_{ij}^{d+1}\Delta_{ij}\Bigg[\left(\partial_{\beta}U_{\lambda}+\partial_{\lambda}U_{\beta}-\frac{2}{d}\delta_{\lambda\beta}\nabla\cdot\mathbf{U}\right)+\frac{d+2}{d}\delta_{\lambda\beta}\nabla\cdot\mathbf{U}\Bigg],$
where the coefficients $B_{k}$ have been defined in Eq. (39). In addition, upon deriving Eq. (B), we have taken into account that by symmetry reasons the only contributions to $P_{\lambda\beta}^{(\Delta\neq 0)c}$ coming from $f_{i}^{(1)}$ only involve the unknowns $\mathcal{C}_{i,\mu\nu}$ and $\mathcal{D}_{i}$. To compute the first term on the right hand side of Eq. (B), we take the leading Sonine approximations to $\boldsymbol{\mathcal{C}}_{i}$ [given by Eq. (94)] and $\mathcal{D}_{i}$:

### v2

@@ Acknowledgements.
We acknowledge financial support from grant no. PID2024-156352NB-I00 funded by MCIU/ AEI/10.13039/501100011033/FEDER, UE and from grant no. GR24022 funded by Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe”. The research of David González Méndez has been supported by the predoctoral fellowship FPU24/01056 from the Spanish Government. AUTHOR DECLARATIONS
@@ Appendix A First-order approximation
$\partial_{t}^{(0)}f_{i}^{(1)}-\sum_{j=1}^{s}\;J_{ij}^{(1)}[f_{i},f_{j}]=-\left(\text{D}_{t}^{(1)}+\mathbf{V}\cdot\nabla+\mathbf{g}\cdot\frac{\partial}{\partial\mathbf{v}}\right)f_{i}^{(0)},$ (138)
where $\text{D}_{t}^{(1)}=\partial_{t}^{(1)}+\mathbf{U}\cdot\nabla$ and the first order contribution $\sum_{j=1}^{s}\;J_{ij}^{(1)}[f_{i},f_{j}]$ to the Enskog collision operator is given by Garzó (2019)
@@
The balance equations at this order provide the actions of the operators $\text{D}_{t}^{(1)}$ over the hydrodynamic fields as
$\text{D}_{t}^{(1)}n_{i}=-n_{i}\nabla\cdot\mathbf{U},\quad\text{D}_{t}^{(1)}\mathbf{U}=-\rho^{-1}\nabla p+\mathbf{g},$ (142)
$\frac{d}{2}n\text{D}_{t}^{(1)}T=-p\nabla\cdot\mathbf{U}-\frac{d}{2}nT\zeta_{U}\nabla\cdot\mathbf{U},$ (143)
@@
$\left(\text{D}_{t}^{(1)}+\mathbf{V}\cdot\nabla+\mathbf{g}\cdot\frac{\partial}{\partial\mathbf{v}}\right)f_{i}^{(0)}=\sum_{j=1}^{s}\Bigg(\mathbf{V}n_{j}\frac{\partial f_{i}^{(0)}}{\partial n_{j}}+\frac{n_{j}}{\rho}\frac{\partial p}{\partial n_{j}}\frac{\partial f_{i}^{(0)}}{\partial\mathbf{V}}\Bigg)\cdot\nabla\ln n_{j}$
@@
$\partial_{t}^{(0)}f_{i}^{(1)}+\left({\cal L}f^{(1)}\right)_{i}=\mathbf{A}_{i}\cdot\nabla\ln T+\sum_{j=1}^{s}\mathbf{B}_{ij}\cdot\nabla\ln n_{j}+C_{i,\lambda\beta}\frac{1}{2}\left(\partial_{\beta}U_{\lambda}+\partial_{\lambda}U_{\beta}-\frac{2}{d}\delta_{\lambda\beta}\nabla\cdot\mathbf{U}\right)+D_{i}\nabla\cdot\mathbf{U},$ (145)
@@
$D_{i}^{\prime}\left(\mathbf{V}\right)=\frac{1}{d}\mathbf{V}\cdot\frac{\partial f_{i}^{(0)}}{\partial\mathbf{V}}+\left(\zeta^{(1,0)}+\frac{2}{d}\frac{p}{nT}\right)T\frac{\partial f_{i}^{(0)}}{\partial T}+\sum_{j=1}^{s}n_{j}\left\{\frac{\partial f_{i}^{(0)}}{\partial n_{j}}+\frac{1}{d}\mathcal{K}_{ij,\beta}\left[\frac{\partial f_{j}^{(0)}}{\partial V_{\beta}}\right]\right\}.$ (157)
@@ Appendix B Collisional contributions to the pressure tensor
$P_{\lambda\beta}^{(\Delta\neq 0)c}$ $=$ $2\sum_{i,j}\chi_{ij}m_{ij}\sigma_{ij}^{d}\Delta_{ij}\int d\mathbf{v}_{1}\int\mathrm{d}\mathbf{v}_{2}\int\mathrm{d}\widehat{\boldsymbol{\sigma}}\,\Theta(\widehat{{\boldsymbol{\sigma}}}\cdot\mathbf{g}_{12})(\widehat{\boldsymbol{\sigma}}\cdot{\bf g}_{12})\widehat{\sigma}_{\lambda}\widehat{\sigma}_{\beta}$
$\times\Bigg[f_{i}^{(1)}(\mathbf{V}_{1})f_{j}^{(0)}(\mathbf{V}_{2})+\frac{1}{2}f_{i}^{(0)}(\mathbf{V}_{1})\frac{\partial f_{j}^{(0)}(\mathbf{V}_{2})}{\partial\mathbf{r}}\cdot{\boldsymbol{\sigma}}_{ij}\Bigg]$
$=$ $\frac{2B_{1}}{d+1}\sum_{i,j}\chi_{ij}m_{ij}\sigma_{ij}^{d}\Delta_{ij}\int\mathrm{d}\mathbf{v}_{1}\int\mathrm{d}\mathbf{v}_{2}g_{12}^{-1}\left(g_{12,\lambda}g_{12,\beta}+g_{12}^{2}\delta_{\lambda\beta}\right)f_{i}^{(1)}(\mathbf{V}_{1})f_{j}^{(0)}(\mathbf{V}_{2})$
$-\frac{B_{2}}{d+2}\sum_{i,j}\chi_{ij}m_{ij}\sigma_{ij}^{d+1}\Delta_{ij}\partial_{\nu}U_{\mu}\int\mathrm{d}\mathbf{v}_{1}\int\mathrm{d}\mathbf{v}_{2}\left(g_{12,\nu}\delta_{\lambda\beta}+g_{12,\lambda}\delta_{\nu\beta}+g_{12,\beta}\delta_{\lambda\nu}\right)f_{i}^{(0)}(\mathbf{V}_{1})\frac{\partial f_{j}^{(0)}(\mathbf{V}_{2})}{\partial V_{2\mu}}$
$=$ $\frac{2B_{1}}{d+1}\sum_{i,j}\chi_{ij}m_{ij}\sigma_{ij}^{d}\Delta_{ij}\int\mathrm{d}\mathbf{v}_{1}\int\mathrm{d}\mathbf{v}_{2}g_{12}^{-1}\left(g_{12,\lambda}g_{12,\beta}+g_{12}^{2}\delta_{\lambda\beta}\right)$
$\times\Bigg[\mathcal{C}_{i,\mu\nu}(\mathbf{V}_{1})\frac{1}{2}\left(\partial_{\mu}U_{\nu}+\partial_{\nu}U_{\mu}-\frac{2}{d}\delta_{\mu\nu}\nabla\cdot\mathbf{U}\right)+\mathcal{D}_{i}\left(\mathbf{V}_{1}\right)\nabla\cdot\mathbf{U}\Bigg]f_{j}^{(0)}(\mathbf{V}_{2})$
$-\frac{B_{2}}{d+2}\sum_{i,j}\chi_{ij}m_{ij}n_{i}n_{j}\sigma_{ij}^{d+1}\Delta_{ij}\Bigg[\left(\partial_{\beta}U_{\lambda}+\partial_{\lambda}U_{\beta}-\frac{2}{d}\delta_{\lambda\beta}\nabla\cdot\mathbf{U}\right)+\frac{d+2}{d}\delta_{\lambda\beta}\nabla\cdot\mathbf{U}\Bigg],$
where the coefficients $B_{k}$ have been defined in Eq. (39). In addition, upon deriving Eq. (B), we have taken into account that by symmetry reasons the only contributions to $P_{\lambda\beta}^{(\Delta\neq 0)c}$ coming from $f_{i}^{(1)}$ only involve the unknowns $\mathcal{C}_{i,\mu\nu}$ and $\mathcal{D}_{i}$. To compute the first term on the right hand side of Eq. (B), we take the leading Sonine approximations to $\boldsymbol{\mathcal{C}}_{i}$ and $\mathcal{D}_{i}$. In the case of $\boldsymbol{\mathcal{C}}_{i}$ is given by Eq. (95) while in the case of $\mathcal{D}_{i}$ is